# RectiCast: Rectifying Distribution Shift in Cascaded Precipitation Nowcasting

### v2

@@ 6 Conclusion
In this paper, we propose RectiCast to address the systematic distribution shift of the deterministic backbone in cascaded precipitation nowcasting methods. By introducing a Rectifier to explicitly rectify the distribution shift and a Generator to focus on modeling local stochasticity, we effectively decouple these two conflicting objectives. Experiments on SEVIR and MeteoNet datasets demonstrate that RectiCast significantly outperforms existing state-of-the-art methods, particularly over longer lead times. Future work will focus on integrating various atmospheric variables to develop a unified precipitation forecasting system.

### v3

@@ 6 Conclusion
In this paper, we propose RectiCast to address the distribution shift of the deterministic backbone in cascaded precipitation nowcasting methods. By introducing a Rectifier to explicitly rectify the distribution shift and a Generator to focus on modeling local stochasticity, we effectively decouple these two conflicting objectives. Experiments on SEVIR and MeteoNet datasets demonstrate that RectiCast significantly outperforms existing state-of-the-art methods, particularly over longer lead times. Future work will focus on integrating various atmospheric variables to develop a unified precipitation forecasting system.